OpenStudy (anonymous):

If y varies inversely as x, and y = 5 as x = 6, find y for the x-value of 10

OpenStudy (lgbasallote):

y varies inversely as x so y = k/x y = 5 when x = 6 5 = k/6 5 x 6 = k 30 = k what is y when x = 10 y = k/x y = 30/10

OpenStudy (lgbasallote):

just tell me if you dont understand soomething